Pulling out of a dive, the pilot of an airplane guides his plane into a vertical circle. At the bottom of the dive, the speed of the airplane is 390 m/s. What is the smallest radius allowable for the vertical circle if the pilot's apparent weight is not to exceed 5.0 times his true weight?
1
Expert's answer
2020-06-17T09:20:36-0400
"N-mg=\frac{mv^2}{r}"

"5mg-mg=\frac{mv^2}{r}\\4g=\frac{v^2}{r}\\4(9.8)=\frac{390^2}{r}"

"r= 3900\ m"